-/1 POINTS SCALC8 4.5.047. Evaluate the definite integral. Need Help? Watch It Talk to a Tutor
Answer #1

So)? when bet dr3= ² . s dx = atat 7:3 7 Z=0 and, when Z=1 x=4 on sulshituting we get $ 157-a de = 5(2-3). + (27dt) - 2 4 1 =